OpenStudy (anonymous):

A country's population in 1990 was 156 million. In 1996 it was 162 million. Estimate the population in 2016 using the exponential growth formula. Round your answer to the nearest million.

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

@aric200

OpenStudy (anonymous):

\[162 = 156(1+x)^{6}\]

OpenStudy (anonymous):

So you need to solve for x first

OpenStudy (aric200):

I just checked this, It seems correct http://openstudy.com/study#/updates/5201dbfbe4b0ebbcb9c6b14d @wolf1728 The population increases156,000,000 to 162,000,000 over 6 years 156,000,000 * (1+r)^6 = 162,000,000 (1+r)^6 = 162,000,000 / 156,000,000 (1+r)^6 = 1.03846153846154 6 * log(1+r) = log(1.03846153846154) log(1+r) = 0.01639041618817 / 6 log(1+r) = 0.0027317360313616 1+r = 1.0063098786004 r = .0063098786004 Testing the value of r 156,000,000 * (1+ .0063098786004)^6 = 1996 population 156,000,000 * 1.03846153846154 = 162,000,000 1996 population r = .0063098786004 Plug r in to find pop in 2016.

OpenStudy (anonymous):

how do I solve for x

OpenStudy (anonymous):

\[\frac{ 162 }{ 156 } = (1+x)^6\]

OpenStudy (anonymous):

26?

OpenStudy (anonymous):

\[1.038461538 = (1+x)^6\] \[1.038461538 - 1 =0.38461538\] \[\sqrt[6]{0.038461538} = 0.0063\]

OpenStudy (anonymous):

could u plz just tell me the answer bc I don't get it

OpenStudy (anonymous):

183.717165048

OpenStudy (anonymous):

that's the answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[2016-1990=26\] \[156(1+0.0063098786004)^{26} = 183.717165048\]

OpenStudy (hwyl):

stop giving out answers. cheating will not be tolerated

OpenStudy (wolf1728):

To solve for x you need to tale logs of both sides 162/156=(1+x)^6

OpenStudy (wolf1728):

TAKE logs

OpenStudy (wolf1728):

I guess jherbo isn't here anymore